OpenStudy (anonymous):

use analytic methods to find any point of discontinuity for the given function. f(x)= (7000x)/(2100-2x) i have no clue what to do, can anyone help me please?

OpenStudy (campbell_st):

the discontinuity occurs when the denominator is equal to zero. you need to solve 21 - 2x = 0 that value will be the discontinuity... or vertical asymptote.

OpenStudy (anonymous):

that doesnt seem right it gives me a fraction of 1/1050

OpenStudy (campbell_st):

opps its 2100 - 2x = 0 that value of x is the discontinuity or vertical asymptote.

OpenStudy (anonymous):

what does "analytical methods" mean? take \[2100-2x=0\] and solve for \(x\)

OpenStudy (anonymous):

oh what @campbell_st said

OpenStudy (anonymous):

that gives me 1050 then what do i do do i do 7000x/1050?

OpenStudy (anonymous):

@satellite73

OpenStudy (campbell_st):

here is a picture of whats happeneing  as x gets close to 1050 the function get close to infinity...